# Creativity and Machine Learning: A Survey

### v4

@@ 4.10.2. Dimensions of Creativity Considered
The authors of (Grace and Maher, 2014) suggest that unexpectedness can be used to compute three dimensions of creativity: novelty, surprise and transformativity. Indeed, novelty is about the possibility of violating the observer’s expectations about the continuity of a domain; if the current model of belief is not applicable for the current artifact, it can be considered as novel. Surprise is instead about the possibility of violating the observer’s expectations about an artifact. Finally, transformativity is about the possibility of violating observer’s expectations about the conceptual space itself (i.e., finding that the rules governing it were not accurate).
@@ 4.10.4. Critical Examination
The unexpectedness measure appears to be able to provide researchers and practitioners with a way to derive novelty and surprise. Notably, it also captures transformativity, clarifying at the same time how simple surprise differs from it, i.e., that surprise is related with expectations about a single artifact, while transformativity is related with expectations about the entire domain. However, it requires to define its conceptual space in terms of explicit rules, which can be violated (and in a way that allows a violation to be detected). In addition, it does not include value in the assessment of machine creativity.
@@ 4.11.1. Overview
The metric proposed by (Maher, 2010) is based on three components: novelty, value and surprise. It relies on the idea that a creativity metric has to be independent not only from the domain, but also from the producer.
@@ 4.11.2. Dimensions of Creativity Considered
The criteria of creativity defined by (Maher, 2010) cover exactly the Boden’s three criteria. In particular, novelty is intended here as a measure of how different the artifact is from known artifacts in its class. Value is quantified considering how the potentially creative artifact compares in utility, performance, or attractiveness to other artifacts in its class. Finally, unexpectedness is defined as the variation from expectation developed for the next new artifact in its class.
@@ 4.11.3. Protocol for Evaluation
Novelty is calculated as the distance between the artifact of interest and the other artifacts in its class. The partition into classes is obtained by means of a clustering algorithm. Surprise is calculated considering whether or not the artifact agrees with the expected next artifact in the pattern extracted from recent artifacts. More specifically, it is calculated as the difference between the expected next artifact and the real next artifact. Such a pattern is predicted by a self-supervised neural network; predictions are refined using reinforcement learning to correct the learned trajectory in case of sequential data. Finally, value is calculated as the weighted sum of the performance variables of the artifact. The weights depend on a co-evolutionary algorithm with a fitness function that can change over time in case the current population of artifacts changes.
@@ 4.11.4. Critical Examination
The method considers all the three Boden’s criteria; it is not linked to a specific domain, or the producer itself; it deals with the evolution of creativity, capturing its volatile nature at the same time. However, in our opinion, it is limited in terms of applicability by the fact that it requires the definition of performance variables (similarly to other approaches based on attribute-value pairs, see Section 4.8.4). Moreover, the setting of the parameters of the clustering algorithms at the basis of this method and the definition of distances among artifacts require human fine-tuning.
@@ 4.12.1. Overview
For the specific case of storytelling, the authors of (Karampiperis et al., 2014) propose a set of computational metrics in order to compute evaluation of novelty, surprise, rarity and recreational effort.
@@ 4.12.2. Dimensions of Creativity Considered
Novelty and surprise are evaluated according to the standard Boden’s definition, while rarity is intended as the presence of rare combinations of properties and recreational effort as the difficulty to achieve a specific result.
@@ 5. Outlook and Conclusion
Even if the field of machine creativity has witnessed an increasing interest and popularity in the recent years, there are still several open challenges. First of all, an interesting direction is the exploration of creativity-oriented objective functions, to directly train models to be creative, or to navigate the induced latent space to find creative solutions. Another open problem is the definition of techniques to explore or transform the space of solutions. A fundamental area is the definition of novel and robust evaluation techniques for both generated and real artifacts. As discussed in Section 4, deep learning might be used as a basis for the definition of metrics for machine creativity. It should be noted that there is also an ongoing debate about the role of human versus machine evaluation (Lamb et al., 2018).
Another promising research direction concerns the machine interpretation of art (Achlioptas et al., 2021). Moreover, machine learning techniques might also be used to investigate psychological dimensions of creativity (Agnoli et al., 2019). There are also foundational questions related to generative deep learning and copyright (Franceschelli and Musolesi, 2022). For example, it is not clear if machine-generated works could be protected by Intellectual Property, and, if they are, who should be the owner of the related rights (Samuelson, 1986). In addition, other problems concerning copyright should be considered, such as if and when training over protected work is permitted (Sobel, 2020). Another important ongoing debate is about authorship and the human role in creative fields in the era of AI [^7] This is one of the ethical dilemmas highlighted by UNESCO in its Preliminary study on the Ethics of Artificial Intelligence..
The models and framework discussed in this work show the remarkable potential of generative learning for machine creativity. We hope that this survey will represent a valuable guide for researchers and practitioners working in this fascinating cross-disciplinary and trans-disciplinary area.

### v6

@@ 4.9.2. Dimensions of Creativity Considered
The authors of (Grace and Maher, 2014) suggest that unexpectedness can be used to compute three dimensions of creativity: novelty, surprise, and transformativity. Indeed, novelty is about the possibility of violating the observer’s expectations about the continuity of a domain; if the current model of belief is not applicable to the current artifact, it can be considered novel. Surprise is instead about the possibility of violating the observer’s expectations about an artifact. Finally, transformativity is about the possibility of violating the observer’s expectations about the conceptual space itself (i.e., finding that the rules governing it were not accurate).
@@ 4.9.4. Critical Examination
The unexpectedness measure appears to be able to provide researchers and practitioners with a way to derive novelty and surprise. Notably, it also captures transformativity, clarifying at the same time how simple surprise differs from it, i.e., that surprise is related to expectations about a single artifact, while transformativity is related to expectations about the entire domain. However, it requires defining its conceptual space in terms of explicit rules, which can be violated (and in a way that allows a violation to be detected). In addition, it does not include value in the assessment of machine creativity.
@@ 4.10.1. Overview
The metric proposed by (Maher, 2010) is based on three components: novelty, value, and surprise. It relies on the idea that a creativity metric has to be independent not only from the domain but also from the producer.
@@ 4.10.2. Dimensions of Creativity Considered
The criteria of creativity defined by (Maher, 2010) cover exactly Boden’s three criteria. In particular, novelty is intended here as a measure of how different the artifact is from known artifacts in its class. Value is quantified by considering how the potentially creative artifact compares in utility, performance, or attractiveness to other artifacts in its class. Finally, unexpectedness is defined as the variation from expectation developed for the next new artifact in its class.
@@ 4.10.3. Protocol for Evaluation
Novelty is calculated as the distance between the artifact of interest and the other artifacts in its class. The partition into classes is obtained by means of a clustering algorithm. Surprise is calculated by considering whether or not the artifact agrees with the expected next artifact in the pattern extracted from recent artifacts. More specifically, it is calculated as the difference between the expected next artifact and the real next artifact. Such a pattern is predicted by a self-supervised neural network; predictions are refined using reinforcement learning to correct the learned trajectory in case of sequential data. Finally, value is calculated as the weighted sum of the performance variables of the artifact. The weights depend on a co-evolutionary algorithm with a fitness function that can change over time in case the current population of artifacts changes.
@@ 4.10.4. Critical Examination
The method considers all three of Boden’s criteria; it is not linked to a specific domain, or the producer itself; it deals with the evolution of creativity, capturing its volatile nature at the same time. However, in our opinion, it is limited in terms of applicability by the fact that it requires the definition of performance variables (similarly to other approaches based on attribute-value pairs, see Section 4.8.4). Moreover, the setting of the parameters of the clustering algorithms at the basis of this method and the definition of distances among artifacts require human fine-tuning.
@@ 4.11.1. Overview
For the specific case of storytelling, the authors of (Karampiperis et al., 2014) propose a set of computational metrics to compute the evaluation of novelty, surprise, rarity, and recreational effort.
@@ 4.11.2. Dimensions of Creativity Considered
Novelty and surprise are evaluated according to the standard Boden’s definition, while rarity is intended as the presence of rare combinations of properties and recreational effort as the difficulty in achieving a specific result.
@@ 5. Outlook and Conclusion
Even if the field of machine creativity has witnessed increasing interest and popularity in recent years, there are still several open challenges. First of all, an interesting direction is the exploration of creativity-oriented objective functions, to directly train models to be creative or to navigate the induced latent space to find creative solutions. Another open problem is the definition of techniques to explore or transform the space of solutions. A fundamental area is the definition of novel and robust evaluation techniques for both generated and real artifacts. As discussed in Section 4, deep learning might be used as a basis for the definition of metrics for machine creativity. It should be noted that there is also an ongoing debate about the role of human versus machine evaluation (Lamb et al., 2018). Another promising research direction concerns the machine interpretation of art (Achlioptas et al., 2021). Moreover, machine learning techniques might also be used to investigate psychological dimensions of creativity (Agnoli et al., 2019). There are also foundational questions related to generative deep learning and copyright (Franceschelli and Musolesi, 2022). For example, it is not clear if machine-generated works could be protected by Intellectual Property, and, if they are, who should be the owner of the related rights (Samuelson, 1985). In addition, other problems concerning copyright should be considered, such as if and when training over protected work is permitted (Sobel, 2020). Another important ongoing debate is about authorship and the human role in creative fields in the era of AI [^5] This is one of the ethical dilemmas highlighted by UNESCO in its Preliminary study on the Ethics of Artificial Intelligence..
The models and framework discussed in this work show the remarkable potential of generative learning for machine creativity. We hope that this survey will represent a valuable guide for researchers and practitioners working in this fascinating area.